What is the percent ionization of a 0.213 M acetic acid solution with a Ka = 1.75×10−5? Round answer to 3 sig figs and do not include the percent sign.


please show work
Answer #1

% jonisation - - X 100 [CH3COOH] from ICE table a can be calculated of acetate ion after dissociation x is concentration of aт” 5 xia®хо-эз - 0-313 хо* * = 3-73 x * 12 | 313xo 6 = 1.43x108 From equilibrium -3 Гcatoo l = = 1.43х10 Letoo 2 х 100 :. 7